OpenStudy (anonymous):

Solve –2x2 +3x – 9 = 0

OpenStudy (deez):

-2x^2 + 3x - 9 = 0 => 2x^2 -3x - 9 =0 (2x-3) (x+3)=0 X=-3, X= 3\2

OpenStudy (anonymous):

this is quadratic equations with complex numbers

OpenStudy (saifoo.khan):

i guess the answers are wrong! :O

OpenStudy (anonymous):

no,i just need it for another prob

OpenStudy (saifoo.khan):

Em, plz write here. the chat isnt working.